OpenStudy (anonymous):

prove by induction that 7^n-4^n-3^n is divisible by 12, when n is a member of positive integers

OpenStudy (anonymous):

you mean when n > 1 ?

OpenStudy (anonymous):

when n>0

OpenStudy (anonymous):

so do u get it? I really don't know how to do this

OpenStudy (anonymous):

ok anyway .. check for n = 1 7^1 - 4^2 - 3^2 = 0 divisible by 12 check for n = 2 : 7^2 - 4^2 - 3^2 = 24 which is divisible by 12 now assume for n=k that (7^k - 4^k - 3^k) / 12 divisible so : no lets check for n+1 : (7^(k+1) - 4^(k+1) - 3^(k+1)) / 12 7*7k/12 -4*4^k/12 - 3*3^k/12 (3+4)*7^k/12 - 4*4^k/12 - 3*3^k/12 3*7^k/12 + 4*7^k/12 - 4*4^k/12 - 3*3^k/12 (7^k - 3^k) /4 + (7^k - 4^k)/3 now for each we have to check that they are divisible (7^k - 3^k)/4 k = 1 true k = 2 true assume k=t divisible (7^t - 3^t)/ 4 k= t+ 1 (7*7^t - 3*3^t )/4 (3*7t + 4*7t - 3*3^t) / 4 7^t + 3*(7^t - 3^t)/4 <- divisible by assumption do the same with : (7^k - 4^k)/3

OpenStudy (anonymous):

hope you are still here and sorry that it took me forever

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

you are welcome in fact it was really challenging, almost gave up im happy to find such questions